Evaluate the triple integral ∭E z dV where E is the solid bounded by the cylinder y^2+z^2=576 and the planes x=0,y=4x and z=0 in the first octant.
1
Expert's answer
2020-11-17T05:33:49-0500

"\iiint_E zdV,"

"E=\{y^2+z^2 = 256, y=4x,x=0,z=0, x>0,y>0,z>0\}"

"z \in[0;\sqrt{256-y^2}]\\
x\in[0;\frac{y}{4}]\\
y\in[0;16]"

"\int_0^{16}dy\int_0^{\frac{y}{4}}dx\int_0^{\sqrt{256-y^2}}zdz = \\
=\int_0^{16}dy\int_0^{\frac{y}{4}}dx(\frac{z^2}{2}|_0^{\sqrt{256-y^2}}) = \\=
\int_0^{16}dx\int_0^{\frac{y}{4}}(\frac{256-y^2}{2})dx = \int_0^{16}(128-\frac{y^2}{2})\frac{y}{4}dy=\\=
\int_0^{16}(32y-\frac{y^3}{8})dy = 16y^2 - \frac{y^4}{32}|_0^{16} = 4096-2048=\\=
2048"